OpenStudy (anonymous):

\[\int\limits [1/(\sec ^2x \csc ^5x)\]

OpenStudy (turingtest):

\[{1\over\sec^2 x\csc^5 x}=\cos^2x\sin^5x=\cos^2x\sin^4x (\sin x)=\cos^2x(1-\cos^2)^2\sin x\]expand and u-sub\[u=\cos x\implies du=-\sin xdx\]

OpenStudy (anonymous):

thanks I got it the answer is (u^3/3)-(2u^5/5)+(u^7/7), am i right?

OpenStudy (turingtest):

I don't know, I haven't checked!

OpenStudy (turingtest):

one sec, brb

OpenStudy (turingtest):

umm, besides missing the minus signs I thin that is right

OpenStudy (turingtest):

remember derivative cos is -sin, so all these u-subs need a -1 in front -(u^3/3)+(2u^5/5)-(u^7/7)

OpenStudy (anonymous):

oh ya i see thanks again

OpenStudy (turingtest):

welcome :D and don't forget +C !!!

OpenStudy (anonymous):

yes thank you